### v2

# Is Complex Query Answering Really Complex?

Abstract
Complex query answering (CQA) on knowledge graphs (KGs) is gaining momentum as a challenging reasoning task. In this paper, we show that the current benchmarks for CQA might not be as complex as we think, as the way they are built distorts our perception of progress in this field. For example, we find that in these benchmarks, most queries (up to 98% for some query types) can be reduced to simpler problems, e.g. link prediction, where only one link needs to be predicted.The performance of state-of-the-art CQA models decreases significantly when such models are evaluated on queries that cannot be reduced to easier types. Thus, we propose a set of more challenging benchmarks composed of queries that require models to reason over multiple hops and better reflect the construction of real-world KGs. In a systematic empirical investigation, the new benchmarks show that current methods leave much to be desired from current CQA methods.
Machine Learning, ICML
1 Introduction
Figure 1: Query answers are not all equally hard when some links can be found in the training data as shown for the 2i1p query $q_{1}$ and fragments of the KG $\mathsf{FB15k237}$, where $r_{1}=\mathsf{distributedVia}^{-1}$, $r_{2}=\mathsf{locatedIn}^{-1}$ and $r_{3}=\mathsf{performedIn}^{-1}$. Its different answers can be obtained by traversing the training graph (continuous line) and predicting the missing links (dotted lines). (Top) Example answer that requires all links to be predicted. (Bottom) Example answers that require only a subset of the links to be predicted and that, therefore, can be reduced to the simpler types 1p, 2p, and 2i (see Sec. 2 and Figure 2).
A crucial challenge in AI and ML is learning to perform complex reasoning, i.e., solving tasks that involve several intermediate steps and sub-goals to be completed. Complex query answering (CQA; Hamilton et al., 2018; Zhang et al., 2021; Arakelyan et al., 2021; Zhu et al., 2022) emerged as a way to measure complex reasoning over external knowledge bases, encoded as knowledge graphs (KGs; Hogan et al., 2021). For instance, to answer the query:
Which actor performed in a movie filmed in ($q_{1}$)
New York City and distributed on Blue Ray?
over a KG such as FreeBase (Bollacker et al., 2008), one would need to first intersect the set of movies found on Blue Ray and the ones shot in New York City, and then link these intermediate candidate answers to another entity, i.e., an actor participating in it. However, the answers computed in this way may not include entities that are unreachable if missing links are present. To deal with the unavoidable incompleteness of real-world KGs, ML methods were developed to solve CQA in the presence of missing links. The current state-of-the-art (SoTA) for CQA are neural query answering models,(Arakelyan et al., 2021; Zhu et al., 2022; Arakelyan et al., 2023; Ren et al., 2023; Galkin et al., 2024b), mapping queries and KGs (i.e., entities and relation names) into a unified latent space. Performance measured on de-facto-standard benchmarks such as $\mathsf{FB15k237}$ (Toutanova & Chen, 2015) and $\mathsf{NELL995}$ (Xiong et al., 2017) suggests impressive progress achieved in recent years on CQA on queries having different structures, and hence posing apparently different levels of difficulty to be answered.
The difficulty of a benchmark relates to the size and structural complexity [^1] Not to be confused with the computational complexity of query answering (Dalvi & Suciu, 2012). of its queries, and several query “types” have been proposed (Ren et al., 2020), each involving a different combination of logical operators—conjunctions, disjunctions, and requiring to traverse a number of missing links that generally increases with the number of logical conditions imposed (Figures 1 and 2). For instance, the query $q_{1}$ is an example of a “2i1p” query type, since it comprises an intersection of two entity sets (2i) followed by a path [^2] Also known as projection in related works (Ren et al., 2023). of length one (1p). In this paper, we argue that the perception of progress on CQA benchmarks has been distorted by implicit assumptions in these benchmarks. We start by noting how both in $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, the vast majority of queries of a complex type simplify to one simpler type, thanks to the fact that to answer them, one can leverage links already appearing in the training data. Figure 1 shows an example of how answers to the query $q_{1}$ on $\mathsf{FB15k237}$– which should require predicting three missing links – require the same effort associated with simpler types involving fewer links. In fact, in $\mathsf{FB15k237}$, the answer “K. Dunst” can be retrieved by predicting just one missing link when leveraging the training links. In this sense, a 2i1p query reduces to a “1p” query, i.e., the much simpler task of link prediction (see Sec. 2). Similarly, the answers “J. Bryant” and “K. Chandler” can be retrieved by predicting two links instead of three. Only the answer “A. Serkis” requires predicting three missing links, and thus follows the intuitive “hardness” associated to the type 2i1p, i.e., it is the hardest when no information is available. As we note here, there is a spectrum of hardness, predicting “K. Dunst” will be easier than answering “K.Chandler”, which in turn is easier than saying “A. Serkis”. However, current CQA benchmarks greatly overrepresent the easy query answers over the others – the “K. Dunst”-like answers are the vast majority – and this inflates performance as SoTA models are prone to memorize training links. In this paper, we analyze this issue and propose a new set of benchmarks that are balanced and cover the full spectrum of hardness in CQA.
Contributions. After revisiting the CQA task (Sec. 2) and highlighting how query types can simplify at test time in the presence of training links (Sec. 3), (C1) we show that up to 99.9% of the test queries can be reduced to simpler queries, with the majority of them (up to 98%) reduced to “one-step” link prediction problems (Sec. 4). This clearly inflates the reported performance. (C2) We re-evaluate previous SoTA approaches (Sec. 5), revealing that neural link predictors rely on memorized information from the training set. Furthermore, we show that the reported hardness of queries involving unions is only apparent. To better understand why performances are inflated, we introduce CQD-Hybrid, a hybrid CQA solver that combines classical graph matching with neural link predictors. (C3) We create new benchmarks for CQA (Sec. 6), $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$ from $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, respectively, where we equally sample all intermediate query types one can consider by accounting for known links in the training data. To these, we build $\mathsf{ICEWS18}$+$\mathsf{H}$ from the temporal KG $\mathsf{ICEWS18}$ (Boschee et al., 2015), where validation and test links are links that have been added to the KG after those in the training, to investigate more realistic sampling schemes to generate the train/validation/test splits. Lastly, we add the query types “4p” (a four-length path) and “4i” (a conjunction of four patterns), making them more challenging.
2 KGs and Complex Query Answering
Knowledge graphs. A KG is a graph-structured knowledge base where information is encoded as relationships between entities. More formally, a KG is a multi-relational graph $G=(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$ is a set of entities, $\mathcal{R}$ is a set of relation names, and $\mathcal{T}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ is a set of links or triples, where each triple $(s,p,o)\in\mathcal{T}$ represents a relationship of type $p$ between a subject $s$ and an object $o$. For instance, in a KG such as FreeBase, the fact that the movie “Spiderman 2” is distributed in Blue Ray is stored in the triple $(\mathsf{Spiderman2},\mathsf{distributedVia},\mathsf{BlueRay})$ or equivalently $(\mathsf{BlueRay},\mathsf{distributedVia}^{-1},\mathsf{Spiderman2})$ where $\mathsf{distributedVia}^{-1}$ is the inverse relation of $\mathsf{distributedVia}$. Figure 1 shows examples of fragments of the KG $\mathsf{FB15k237}$.
The aim of CQA is to retrieve a set of answers to a logical query $q$ that poses conditions over entities and relation types in a KG. Following Arakelyan et al. (2021), we consider the problem of answering logical queries with a single target variable ($T$), a set of constants including anchor entities ($a_{1},\ldots,a_{k}\in\mathcal{E}$), given relation names ($r_{1},\ldots,r_{n}\in\mathcal{R}$), and first-order logical operations that include conjunction $\wedge$, disjunction $\vee$, negation $\neg$ and existential quantification $\exists$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Query structures we consider, adapted from Ren & Leskovec (2020) and including path (p), intersection (i), union (u) structures. Sec. 2 for their logical formulation. See App. B for additional query structures including negation.
Different queries are categorized into different types based on the structure of their corresponding logical sentence (Xiong et al., 2017). This taxonomy should imply that queries of the same type share the same “hardness”, i.e., the level of difficulty to be answered, and different types correspond to tasks that map to more or less complex reasoning tasks. As we will show in the next sections, this is not the case, and to fully grasp the actual hardness to answer a query one needs to account for the “leakage” of intermediate links. We now review the most common query types used in CQA. The simplest CQA task is link prediction, i.e., answering a query of the form:
?$T:(a_{1},r_{1},T),$ (1p)
that is, given an entity $a_{1}$ (e.g., $\mathsf{NYC}$) and a relation name $r_{1}$ (e.g., $\mathsf{locatedIn}^{-1}$), find the entity that when substituted to $T$ correctly matches the link in the KG (e.g., $\mathsf{Spiderman2}$, see Figure 1). Instead of matching a single link, more complex queries involve matching sub-graphs in a KG (see Figure 2). Xiong et al. (2017) extend 1p queries, and ask questions involving sequential paths made of two or three links, i.e.,
$\exists V_{1}.(a_{1},r_{1},V_{1})\land(V_{1},r_{2},T),$ (2p)
$\exists V_{1},V_{2}.(a_{1},r_{1},V_{1})\land(V_{1},r_{2},V_{2})\land(V_{2},r_{3},T),$ (3p)
where $V_{1},V_{2}$ denote variables to be grounded into entities appearing in the path. Moreover, multiple ground entities can participate in a conjunction, e.g., in queries such as:
?$T:(a_{1},r_{1},T)\land(a_{2},r_{2},T),$ (2i)
?$T:(a_{1},r_{1},T)\land(a_{2},r_{2},T)\land(a_{3},r_{3},T),$ (3i)
which represent the intersection of the target entity sets defined over two (2i) or three (3i) links. Path and intersection structures can be combined into more complex queries: for example, the natural language expression for query $q_{1}$ can be formalized as the formula
?$T:\exists V_{1}.(a_{1},r_{1},V_{1})\land(a_{2},r_{2},V_{1})\land(V_{1},r_{3},T),$ (2i1p)
involving one intersection followed by a length-one path. [^3] In Ren et al. (2020) this type was referred to as “ip”. We explicitly mention the number of steps involved in a path or conjunction, as this is a factor of complexity. Analogously, “pi”, and “up” queries from Ren et al. (2020) are now 1p2i, and 2u1p. See also the example in Figure 1. By inverting the order of operations, we obtain the query type “1p2i”:
?$T:\exists V_{1}.(a_{1},r_{1},V_{1})\land(V_{1},r_{2},T)\land(a_{2},r_{3},T).$ (1p2i)
Similarly to introducing conjunctions, we can consider disjunctions in queries, realizing the union query types which can be answered by matching one link or the other, such as
?$T:(a_{1},r_{1},T)\lor(a_{2},r_{2},T),$ (2u)
or by combining the union with the previous query types, e.g., obtained by combining 2u and 1p:
?$T:\exists V_{1}.((a_{1},r_{1},V_{1})\lor(a_{2},r_{2},V_{1}))\land(V_{1},r_{2},T).$ (2u1p)
Note that despite their dissimilar syntaxes and the associated sub-graphs (Figure 2), the query type 2u should be as difficult as 1p, as to answer the first it suffices to match a single link correctly. Similarly, 2u1p is as complex as 2p. The fact that 2u and 2u1p are reported to be harder to solve in practice than 1p and 2p (Ren et al., 2020) is due to the way standard benchmarks are built, which we discuss in Sec. 4, and the way in which CQA is evaluated, discussed next. App. B details further query types, including negation.
Standard evaluation. Given a KG $G$ and a query $q$, answering $q$ boils down to a graph matching problem (Hogan, 2020) if we assume that all the meaningful links are already in $G$. Instead, if $G$ is incomplete, we will need to predict missing links while answering $q$. Many ML approaches to CQA, reviewed in Sec. 5, therefore assume a distribution over possible links (Loconte et al., 2023), requiring probabilistic reasoning. To evaluate them, standard benchmarks such as $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ artificially divide $G$ into $G_{\operatorname{train}}$ and $G_{\operatorname{test}}$, treating the triples in the latter as missing links. This splitting process is done uniformly at random, a procedural choice that can alter the measured performance, as we discuss next. CQA is generally treated as a ranking task, counting how many true candidate answers to a query are ranked higher than non-answer ones. Let the rank of a query answer (QA) pair $(q,t)$ be $\operatorname{rank}(q,t)$, the performance for each query type is calculated as the mean reciprocal rank (MRR), i.e.:
$|\mathcal{Q}|^{-1}\sum\nolimits_{q\in\mathcal{Q},t\in\mathcal{E}_{q}}\left|\mathcal{E}_{q}\right|^{-1}\operatorname{rank}(q,t)^{-1},$ (1)
where $\mathcal{Q}$ denotes the set of test queries of the considered type, and $\mathcal{E}_{q}$ is the set of candidate answer entities for each query $q\in\mathcal{Q}$. This average, across queries and answers, assumes that every QA pair having the same query type is equivalently hard, which is not the case. In fact, we show that certain QA pairs can be easier to predict if some links were already seen by the model during training (Figure 1) and that the distribution of the query answer pairs in the existing benchmarks is very skewed towards those involving a single missing link (Table 1). Thus, computing Equation 1 without understanding what the benchmark distribution is distorts the perception of performance gains.
3 What is the real “Hardness” of CQA in incomplete KGs?
As discussed in the previous section, the perceived complexity of a query is related to the graph structure associated to its query type (Figure 2): queries containing more hops/existentially-quantified variables are more challenging, e.g., a 3p query is harder than a 2p query. In this section, we give an alternative perspective on the difficulty of answering queries that take into account the information coming from the training data. We argue that predicting links that are truly missing, i.e., not accessible to a learned model, is actually what makes a query “hard”. To do so, we formalize the notion of a reasoning tree for a QA pair, and then define how we determine the practical hardness of a QA pair.
Table 1: The great majority of QA pairs are of the much easier partial-inference type (non-diagonal), rather than of the harder full-inference type (diagonal). For each query type (as rows) we show the percentage of the QA pairs that can be reduced to an easier query type (as columns) for datasets $\mathsf{FB15k237}$ and $\mathsf{NELL995}$. Most of the complex queries can be reduced to simple link prediction queries (i.e., 1p). We denote as ‘-’ those reductions that are not possible given the sub-graph structure induced by the query type.
$\mathsf{FB15k237}$ $\mathsf{NELL995}$
1p 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p 1p 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p
1p 100 - - - - - - - - 100 - - - - - - - -
2p 98.1 1.9 - - - - - - - 97.6 2.4 - - - - - - -
3p 97.2 2.7 0.1 - - - - - - 95.6 4.3 0.1 - - - - - -
2i 96.0 - - 4.0 - - - - - 94.0 - - 6.0 - - - - -
3i 91.6 - - 8.2 0.2 - - - - 87.4 - - 12.1 0.5 - - - -
1p2i 86.8 1.0 - 12.0 - 0.2 - - - 49.5 0.6 - 49.0 - 0.9 - - -
2i1p 96.7 1.8 - 1.4 - - 0.1 - - 96.2 2.4 - 1.2 - - 0.2 - -
2u 0.0 - - - - - - 100 - 0.0 - - - - - - 100 -
2u1p 98.3 0.0 - - - - - 1.6 0.1 98.5 0.0 - - - - - 1.4 0.1
Given a query $q$ and every answer set of candidate answers $t\in\mathcal{E}_{q}$, we define the reasoning tree of each QA pair $(q,t)$, as the directed acyclic graph starting from the anchor entities of $q$ to the target entity $t$, whose relational structure matches the query graph. Figure 1 provides examples of different reasoning trees for four different answers to the same query. There, we highlight whether a link belongs to $G_{\text{train}}$ or not, i.e., it is a missing link. We assess the hardness of each answer $t\in\mathcal{E}_{q}$, by analyzing the composition of the reasoning tree required to predict $t$. As to answering a $(q,t)$ pair, multiple reasoning trees are possible, so we select the one with the smallest number of missing links and the fewest number of hops. A $(q,t)$ pair is trivial if the answer $t$ can be entirely retrieved from $G_{\text{train}}$, i.e., there exists at least one reasoning tree where all the links in it are seen during training. This type of answer does not need probabilistic inference and hence is filtered out from current CQA benchmarks (Ren & Leskovec, 2020), which only consider non-trivial $(q,t)$ pairs, which we call inference-based pairs. However, inference-based pairs do not need all links in their reasoning tree to be predicted as, by definition, it is sufficient that at least one link in the tree is missing. Therefore, we define a $(q,t)$ pair to be a partial-inference pair if its reasoning tree contains at least one link in $G_{\text{train}}$ and at least one link present in $G_{\text{test}}$. Alternatively, a QA pair whose reasoning tree contains only links present in $G_{\text{test}}$ is called a full-inference pair. We claim that inference-based queries encompass a spectrum of hardness, with full-inference queries being the hardest and queries that have only one truly missing link (1p) the easiest. However, current benchmarks flatten this complexity and highlight only performance on the easiest.
To predict a partial-inference pair $(q,t)$, a ML model that has explicit access to $G_{\operatorname{train}}$ has to solve a simpler task to answer $q$, as only a subset of the links in the reasoning tree are missing and need to be predicted. As such, the $(q,t)$ pair can be simplified to $(q^{\prime},t)$ pair, where the query $q^{\prime}$ is of an easier query type than $q$. Figure 1 shows one example of how a QA pairs of type 2i1p from $\mathsf{FB15k237}$ are reduced in practice to the simpler types 1p, 2i and 2p. This is exploited by hybrid solvers such as QTO (Bai et al., 2023) and our CQD-Hybrid (Sec. 5). Note that this advantage applies also to ML models that have implicit access to $G_{\operatorname{train}}$, e.g., by having memorized the triples during training, a common phenomenon for many neural link predictors (Nickel et al., 2014). We confirm this in Sec. 5.
4 How Many “Easy” Queries in Current CQA Benchmarks?
In this section, we systematically analyze the practical hardness of queries from the current CQA benchmarks and answer the following research question: (RQ1) What is the proportion of QA pairs that can be classified as full-inference rather than partial-inference and how easy are the latter? To this end, we consider the CQA benchmarks generated from $\mathsf{FB15k237}$ (based on Freebase) and $\mathsf{NELL995}$ (based on NELL systems (Carlson et al., 2010)) as they are the most used to evaluate SoTA methods for CQA (Ren et al., 2020; Ren & Leskovec, 2020; Arakelyan et al., 2021; Zhang et al., 2021; Zhu et al., 2022; Arakelyan et al., 2023). [^4] We do not consider $\mathsf{FB15k}$ (Bordes et al., 2013) as it suffers from data leakage (Toutanova & Chen, 2015).
Complex queries can be reduced to much simpler types. We group the testing QA pairs $(q,t)$ into query types (Sec. 2), and we further split them based on whether they can be reduced to simpler types after observing the training links in $G_{\operatorname{train}}$ (Sec. 3). Table 1 shows that for both $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ the vast majority of $(q,t)$ pairs can be reduced to simpler types. For $\mathsf{FB15k237}$, 86.8% to 98.3% of $(q,t)$ pairs can be reduced to 1p queries, while only 0.1% to 4% require full inference. Similarly, for $\mathsf{NELL995}$, 49.5% to 98.5% of $(q,t)$ pairs map to 1p queries, and only 0.1% to 6% to full inference. For instance, 96.7% of 2i1p $(q,t)$ pairs in $\mathsf{FB15k237}$ can be reduced to 1p queries, 1.8% to 2p, and 1.4% to 2i. However, only 0.1% of these pairs cannot be reduced to any other $(q^{\prime},t)$ pair, i.e., they require full inference in order to be predicted. The only exceptions to this trend are $(q,t)$ pairs where $q$ has a 1p or 2u structure which, by definition, only require the prediction of a single link and therefore cannot be reduced by any other query type. We conduct the same analysis for queries that include negation, considering only the positive reasoning tree (see Sec. A.1 for details). Table A.1 presents results analogous to those in Table 1, showing that the vast majority of $(q,t)$ pairs including negation have training links in their reasoning tree. For example, 95.4% of 3in $(q,t)$ pairs in $\mathsf{FB15k237}$ have training links in their reasoning tree.
Is this realistic? This skewed distribution is the result of implicit assumptions behind how these benchmarks have been created. We argue they are arbitrary and not realistic. First, we note that the procedure to generate queries assumes that triples are independent, a simplifying assumption that does not hold in practice. Second, as triples are sampled independently, the quantity of QA pairs that can be reduced to 1p depends on how many links are assumed to be missing. This is given by the ratio of $|G_{\operatorname{train}}|$ and $|G_{\operatorname{test}}|$. Note that however the train/test split is arbitrary and has been inherited from the benchmarks for link prediction, where the task was rightfully to predict a single link.
Non-existing links for union queries. Finally, we discovered that there are non-existing links, i.e., links that are not in the original KG $G$ and hence neither in $G_{\operatorname{train}}$ or $G_{\operatorname{test}}$, in both $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ in the reasoning trees of queries involving unions, i.e., the 2u and 2u1p types. Figure A.1 shows some examples. These links violate our definitions of inference ($q$, $t$) pairs, hence, we filter them out, and report only filtered $(q,t)$ pairs in Table 1. More crucially, these non-existing links can alter the performance of solvers for 2u and 2u1p types, as we discuss next.
Takeaway 1. We discourage using current CQA benchmarks as they essentially reflect only the performance of predicting answers where only one link is truly missing (for both hybrid and neural solvers).
3p 2i1p
method ovr 1p 2p 3p 2i 3i 1p2i 2i1p ovr 1p 2p 3p 2i 3i 1p2i 2i1p
GNN-QE 11.8 12.0 4.4 4.8 - - - - 18.9 19.2 8.2 - 6.2 - - 3.4
ULTRAQ 8.9 9.0 4.6 4.4 - - - - 18.6 18.9 8.5 - 5.1 - - 8.1
ConE 11.0 11.0 5.4 2.6 - - - - 14.0 13.8 9.7 - 12.8 - - 5.6
CQD 7.8 7.8 3.4 1.8 - - - - 20.7 21.0 10.5 - 6.7 - - 7.6
CQD-Hybrid 11.0 11.1 3.6 4.4 - - - - 24.0 24.6 10.0 - 4.6 - - 7.5
QTO 15.6 15.8 4.5 5.0 - - - - 24.7 25.3 10.3 - 8.6 - - 8.1
Table 2: SoTA performance drops significantly on full-inference QA pairs as shown for the MRR for 3p and 2i1p queries on $\mathsf{FB15k237}$. Compare the MRR achieved on all available testing queries (“ovr”) with that achieved on the 1p reductions (i.e., when only a single link is truly missing): they are extremely close.
5 What do SoTA Models Perceive as Hard?
In this section, we re-evaluate SoTA methods for CQA and answer to the following question: (RQ2) How do they perform on QA pairs that are placed differently in the hardness spectrum? To substantiate that memorizing training links helps in answering partial-inference queries but not full-inference ones (Sec. 3), we evaluate hybrid solvers who have access to $G_{\operatorname{train}}$.
Table 3: Queries of type 2u are less hard than previously thought in terms of MRR, with our filtered 2u queries having comparable performance to 1p, as discussed in Sec. 2, in contrast with the results of the 2u queries originally used (2u all).
$\mathsf{FB15k237}$ $\mathsf{NELL995}$
method 1p 2u all 2u filtered 1p 2u all 2u filtered
GNN-QE 42.8 16.2 40.7 53.6 15.4 34.8
ULTRAQ 40.6 13.2 33.6 38.9 10.2 21.3
CQD 46.7 17.6 32.7 60.4 19.9 35.9
ConE 41.8 14.9 29.9 60.0 14.9 28.2
Neural CQA methods. Over the years, a significant number of neural models have been proposed for solving the CQA task, see App. C for an overview. Among purely-neural models, we consider four representative approaches that significantly differ in their methodological designs and yield SoTA results compared to other models in their class: (1) Cone Embeddings (ConE; Zhang et al., 2021) is a geometry-based complex query answering model where logical conjunctions and disjunctions are represented as intersections and union of cones. (2) Graph Neural Network Query Executor (GNN-QE; Zhu et al., 2022) decomposes a complex first-order logic query into projections over fuzzy sets. (3) Continuous Query Decomposition (CQD; Arakelyan et al., 2021, 2023), reduces the CQA task to the problem of finding the most likely variable assignment, where the likelihood of each link (1p query) is assessed by a neural link predictor, and logical connectives are relaxed via fuzzy logic operators. (4) UltraQuery (ULTRAQ; Galkin et al., 2024b) is a foundation model for CQA inspired by Galkin et al. (2024a) where links and logical operations are represented by vocabulary-independent functions which can generalize to new entities and relation types in any KG.
Hybrid solvers. So far, we have assessed that current benchmarks are skewed towards easier query types (Table 1) and thus the perceived progress of current SoTA CQA methods boils down to their performance on 1p queries (Table 2). To have an undistorted view of this progress, in the next section, we will devise a benchmark that rebalances all partial-inference as well as full-inference queries types. We argue that in a real-world scenario one has to perform reasoning over both existing links and in the presence of missing ones, hence discarding the information in $G_{\operatorname{train}}$ is wasteful. This, in addition to the fact that purely-neural SoTA CQA methods might implicitly exploit this aspect at test time if they are able to memorize well the entire training KG, pushes us to evaluate also hybrid solvers.
Two remarkable examples are QTO (Bai et al., 2023) and FIT (Yin et al., 2024). They explicitly retrieve existing links from $G_{\operatorname{train}}$ when answering a query at test time. Inspired by them, and to evaluate our hypothesis, we create a light-weight hybrid solver, named CQD-Hybrid, that is a variant of CQD that uses a pre-trained link predictor to compute scores for the answer entities of 1p queries only, denoting the unnormalized probability of that single link to exist (Loconte et al., 2023). Then, assignments to existentially quantified variables in a complex query are greedily obtained by maximizing the combined score of the links, computed by replacing logical operators in the query with fuzzy ones (van Krieken et al., 2022). A complete assignments to the variables (and hence to the target variable $T$), gives us an answer to the query. In our CQD-Hybrid, we assign the maximum score to those links that are observed in $G_{\operatorname{train}}$. That is, training links will have a higher score than the one that the link predictor would output, hence effectively steering the mentioned greedy procedure at test time. This is the only minimal change we apply to CQD to test if its performance can depend on memorizing training triples. Note that this makes CQD-Hybrid different from QTO and FIT, which involve more sophisticated steps such as score calibration and a forward/backward update stage, that can boost performance. In our experiment we will only consider QTO, as for the set of queries we considered QTO and FIT are equivalent (Yin et al., 2024). We report hyperparameters and implementation details of CQD-Hybrid in Sec. F.1.
All SoTA methods perform significantly worse on full-inference queries. We re-evaluate the SoTA methods mentioned above and stratify their performances based on the type of partial- and full-inference queries. Where available, we reused the pre-trained models for their evaluation. Otherwise, we re-produced the SoTA results using the hyperparameters provided by the authors, listed in Sec. F.1.
Table 2 presents a portion of the results in terms of MRR for the testing $(q,t)$ pairs in the existing benchmarks $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, along with the different $(q^{\prime},t)$ pairs they can be reduced to when observing the training KG, grouped by each query type. There, the performance of all SoTA methods consistently drops for query answer pairs that have a higher number of missing links in their reasoning tree. Furthermore, Tables A.5 and A.6 compare the MRR computed on all the available queries originally used in the benchmarks (column “all”), and the scores on those query types that they can be reduced to. There is a high similarity in MRR between the columns “all” and “1p” that is evidence that the good performances of CQA methods are explained by their link prediction performances, as a very high percentage of queries in fact are reduced to 1p queries (see Table 1). A similar conclusion can be drawn for queries including negation, as shown in Table A.2, where the MRR results on partial inference $(q,t)$ pairs are very similar to the overall, and the performance of all SoTA methods consistently drops on full-inference $(q,t)$ pairs.
Note that this happens consistently also for hybrid solvers such as QTO and our CQD-Hybrid: they have a harder time solving QA pairs with more than one missing link. While their performance was consistently surpassing purely-neural solvers, and CQD-Hybrid consistently boosted the performance of CQD, their poor performance on harder QA pairs is not reflected in the overall average score that is usually reported on papers, see Table 2. Moreover, in Table A.3 we highlight that our baseline CQD-hybrid, is able to surpass existing non-hybrid SoTA, which is evidence that a pre-trained neural link predition plus memorization of training links result in SoTA performances for $\mathsf{FB15k237}$ and $\mathsf{NELL995}$. The complete stratified comparison for all solvers is reported in the appendix in Tables A.5 and A.6. However, in a few cases it happens that full-inference results are higher than some partial-inference; in such cases, the reason is to be found in the fact that full-inference $(q,t)$ pairs are very scarce (see Table 1). [^5] We rule out that this is due to the influence of the number of existing entities bounding to existential quantified variables. For example we report such an analysis for CQD in Figures A.3 and D.2. Moreover, we found that no anchor entities nor relation names is predominant in both benchmarks, as shown in Table A.7. This motivates us to create a fully-balanced benchmark in the next section.
Non-existing links for union queries give a false sense of hardness. We remark that for union queries, we filter query answer pairs, as discussed in Sec. 4, removing those with non-existing links in their reasoning trees. In Table 3, we show that if we do not do that, and consider also the pairs with non-existing links (denoted as “2u all”), MRR performance greatly drops. In this way, we reproduce the low performance for 2u queries that the original SoTA baselines reported in their papers. However, as Table 3 shows, this is just an artifact due to including non-filtered pairs while computing Equation 1. With our filtered pairs, instead, we recover the performance of 1p queries, as expected (Sec. 2). Similar conclusions can be drawn for other union queries as reported in Table A.4.
Takeaway 2. $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ are not suitable to precisely assess the capability of CQA methods to answer complex queries, as most of their QA pairs can be predicted by just predicting a single missing links. This results in highly inflated performance that distorts the perception of progress in the field and pushes the community to improve the performance only on the easiest QA pairs.
Figure 3: MRR of SoTA on the new benchmarks is significantly lower than the old ones as reported for $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ (in blue) versus our new $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$ (in orange) for different query types and SoTA models.
6 New Benchmarks for Undistorted CQA Performance
In this section, we aim to answer the following question: (RQ3) How can we construct a set of CQA benchmarks that let us measure the performance on each query type while not distorting the aggregated average? To do so, we present a new set of CQA benchmarks based on the previously used $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, as well as a novel one we design out of temporal knowledge graph $\mathsf{ICEWS18}$ to question the current way to split the original KG $G$ into $G_{\operatorname{train}}$ and $G_{\operatorname{test}}$ (see our discussion of how sampling triples independently is arbitrary in Sec. 4). Then, we evaluate current SoTA methods on these new benchmarks to establish a rigorous baseline for future works.
Building new CQA benchmarks. We generate our benchmarks to comprise both full-inference and partial-inference QA pairs, but we modify the algorithm of (Ren et al., 2020) to ensure that no training links are present in the reasoning trees of full-inference pairs and we balance each query type a QA pair can be reduced to by sampling 10,000 QA pairs each. For example, for 3p queries, we sample 30.000 QA pairs, 10.000 for 3p that can be reduced to 1p, 10.000 for those reduceable to 2p, and 10.000 for full-inference QA pairs (3p QA pairs non-reduceable to any other type). Moreover, to raise the bar of “complexity” in CQA, we introduce two query types that, in their full-inference versions, are harder than simpler types by design. Specifically, we design “4p” queries for a sequential path made of four links and “4i”, which represents the intersection of the target entity sets defined over four links. For their logical formulation, see App. E
We do so for the query types shown in Figures 2 and B.1, and name the harder versions of $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$. Only exceptions are 1p, 2in, 2nu1p those only being of type full-inference. See Secs. A.1 and B for details about queries including negation, and Sec. G.2 for details about query generation.
Non-uniform at random splits. Recall from Sec. 4 that $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ have been created by assuming triple independence and by inheriting the train/test split from their link prediction counterparts. For $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$, we removed the triple independence assumption as we sample the same number of QA pairs per type, but we keep the existing data splits of $G_{\operatorname{train}}$ and $G_{\operatorname{test}}$, for retro-compatibility. However, to evaluate the impact of this artificial splitting process, we adopt a more realistic one for $\mathsf{ICEWS18}$+$\mathsf{H}$, where $G_{\operatorname{train}}$ contains only “past” links, and we might want to predict the future links contained in $G_{\operatorname{test}}$. To this end, we leverage the temporal information in $\mathsf{ICEWS18}$ by (1) ordering the links based on their timestamp; (2) removing the temporal information, thus obtaining regular triples; and (3) selecting the train set to be the first temporally-ordered 80% of triples, the valid the next 10%, and the remaining to be the test split. If the same fact appears with multiple timestamps, we retain only the link with the earliest timestamp. For detailed statistics about the splits, please refer to Sec. G.1. By doing so, we create a much more challenging benchmark even for easy 1p queries.
How do SoTA methods fare on our new benchmarks? For $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$ we re-used the pre-trained models used for the experiments in Sec. 4, while for the newly created $\mathsf{ICEWS18}$+$\mathsf{H}$ we trained GNN-QE, ConE, and the link predictor used in CQD, CQD-Hybrid and QTO from scratch. As this is not necessary for ULTRAQ, being a zero-shot neural link prediction applicable to any KG, we re-used the checkpoint provided by the authors. Hyperparameters of all models are in Sec. F.2.
Figure 3 reports the MRR on the new benchmarks versus the old benchmarks, where the MRR of $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ (in blue) is much higher than the one of $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$ (in orange) for all methods for all query types. Considering how MRR can drop for more than 30 points, e.g., for 3i queries, this highlights once more how distorted the perception of progress in the field can be.
Table 4 shows the results of the selected baselines on the new benchmarks, for all query types in Figure 2. For stratified results on $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$ see Tables D.2, D.3 and D.4, respectively. The surprising result from Table 4 is that no model obtains SoTA performance on most query type, which was not the case for the old benchmarks, where QTO obtained SoTA performance for most query types (see Tables A.6 and A.5). This is further evidence that current SoTA methods mainly focused on improving the performance on the easiest QA pairs and likely overfit the benchmarks. Instead, with our new benchmarks, this will not be possible in the same way, as we balanced the amount of partial- and full-inference QA pairs per type.
Similarly, for queries including negation, the overall MRR greatly drops, and no existing model obtains SoTA for every query type, as shown in Table D.5. Furthermore, we highlight how full-inference 2u queries have similar performance than 1p, similar to how full-inference 2u1p are expected to be as hard as 2p, as discussed in Sec. 3. Finally, we remark that $\mathsf{ICEWS18}$+$\mathsf{H}$ is much harder than $\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{FB15k237}$+$\mathsf{H}$, across all query types, even 1p, thus raising the bar for neural link predictors as well. This confirms our hypothesis that the sampling method of the KG splits plays a big role in determining the hardness of the benchmark.
Table 4: There is no clear SoTA method for the new benchmarks. As shown in Figure 3, the MRR on the new benchmarks is significantly lower than the old ones. For example, for 3i queries on $\mathsf{FB15k237}$+$\mathsf{H}$, QTO has an MRR of 10.1, while for $\mathsf{FB15k237}$ it was 54.6. For the stratified analysis see Tables D.2, D.3 and D.4. For results on queries involving negation see Table D.5.
Model 1p 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p 4p 4i
$\mathsf{FB15k237}$+$\mathsf{H}$ GNN-QE 42.8 5.2 4.0 6.0 8.8 5.6 9.9 32.5 10.0 4.3 20.0
ULTRAQ 40.6 4.5 3.5 5.2 7.2 5.3 10.1 29.4 8.3 3.8 16.4
CQD 46.7 4.4 2.4 11.3 12.8 6.0 11.5 40.1 10.6 1.1 23.8
CQD-Hybrid 46.7 4.8 3.1 6.0 8.6 5.5 12.9 42.2 12.0 2.4 17.4
ConE 41.8 4.6 3.9 9.1 10.3 3.8 7.9 22.8 6.0 3.5 20.3
QTO 46.7 4.9 3.7 8.7 10.1 6.1 13.5 30.6 11.2 3.9 20.2
$\mathsf{NELL995}$+$\mathsf{H}$ GNN-QE 53.6 8.0 6.0 10.7 13.3 16.0 13.5 47.5 9.8 4.7 19.4
ULTRAQ 38.9 6.1 4.1 7.9 10.2 15.8 9.3 28.1 9.5 4.2 15.6
CQD 60.4 9.6 4.2 18.5 19.6 18.9 22.6 46.3 18.5 2.0 24.8
CQD-Hybrid 60.4 9.0 6.1 12.1 14.4 17.4 21.2 46.4 19.3 3.5 20.4
ConE 53.1 7.9 6.7 21.8 23.6 14.9 11.8 39.9 8.8 5.2 27.6
QTO 60.3 9.8 8.0 14.6 15.8 17.6 21.1 49.1 18.9 7.0 20.9
$\mathsf{ICEWS18}$+$\mathsf{H}$ GNN-QE 12.2 0.9 0.5 16.1 26.5 19.1 3.5 7.6 1.1 0.4 34.0
ULTRAQ 6.3 1.2 1.2 7.0 11.7 8.8 1.3 3.3 1.2 0.8 15.9
CQD 16.6 2.5 1.5 13.0 19.5 17.1 6.7 6.8 5.9 1.1 24.0
CQD-Hybrid 16.6 2.6 1.5 15.0 25.5 17.5 5.8 6.8 5.6 0.9 33.2
ConE 3.5 0.9 0.9 1.2 0.5 1.2 1.6 1.1 0.9 0.6 0.3
QTO 16.6 2.6 1.4 15.7 25.0 18.4 6.2 6.7 4.9 1.1 31.5
Takeaway 3. Our benchmarks have a balanced amount of partial-inference and full-inference QA pairs that depends on the query type, allowing to measure CQA performance while not distorting the aggregate performance across all QA pairs with different hardness. Additionally, $\mathsf{ICEWS18}$+$\mathsf{H}$ highlights that more realistic sampling strategies are more challenging for current SoTA.
7 Conclusion
In this paper, we revisit CQA on KGs and reveal that the “good” performance of SoTA approaches predominantly comes from answers that can be reduced to easier types (Table 2), the vast majority of which boiling down to single link prediction (Table 1). We perform such an analysis by dissecting the arbitrary assumptions that were used to create these datasets and by highlighting how current neural and hybrid solvers can exploit (different) forms of triple memorization to make complex queries much easier. We confirm this by reporting their performance in a stratified analysis and by proposing our hybrid solver, CQD-Hybrid, which, while being a simple extension of an old method like CQD, can be very competitive against other SoTA models. This reveals that $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ are not suitable to precisely assess the capability of CQA methods to answer complex queries.
We then created a set of new benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$$\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{ICEWS18}$+$\mathsf{H}$ that comprise a balanced amount of full-inference and partial-inference QA pairs. We consider them to be a stepping stone towards benchmarking truly complex reasoning with ML models. However, both the old and our new benchmarks only consider queries with bindings of single target variables. While this reflects an important class of queries on real-world KGs, many real-world queries require bindings to multiple target variables (i.e. answer tuples). We plan to extend our current study of the “real hardness” of CQA benchmarks to other popular settings in neural query answering, such as inductive scenarios (Galkin et al., 2022) where some entities and/or relations are unseen during the test stage.
Impact Statement
This work analyses current CQA benchmarks for knowledge graphs and shows that the way benchmarks are created inflates perceived progress in the field. By demonstrating that most benchmark queries can be reduced to simpler problems and developing new, more balanced benchmarks, this work helps establish more accurate measures of model performance in CQA tasks. The findings have important implications for how the research community evaluates and develops CQA systems, potentially leading to more robust and capable models to better handle truly complex queries.
Author Contributions
AV, CG, and LL conceived the initial idea for which existing benchmarks were not hard enough for complex query answering. AV, CG, DH, and LL had the intuition that a simple hybrid solver could obtain SoTA results on the old benchmarks. BX, and CG found the issue with union queries. CG developed the necessary code, ran all the experiments, and plotted the figures, with the following exceptions: BX ran all the experiments involving ConE, LL contributed to the generation of the new benchmarks, specifically for the 4p and 4i queries, PM wrote the script for generating the KG splits of ICEWS18, AV drew Fig.1 and 2. BX, and CG wrote the initial draft of the paper, while AV shaped the story line of the paper and helped with the writing. All authors critically revised the paper. AV, PM, and SS supervised all the phases of the project and gave feedback.
Acknowledgements
AV was supported by the “UNREAL: Unified Reasoning Layer for Trustworthy ML” project (EP/Y023838/1) selected by the ERC and funded by UKRI EPSRC. PM was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP. CG was funded by SMARTY (grant no. 101140087). BX was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-1574 – 471687386. DH was supported by the Deutsche Forschungsgemeinschaft (DFG): Germany’s Excellence Strategy – EXC 2120/1, GA 390831618 (RP20).
Appendix A Additional analysis on the old benchmarks
A.1 Influence of existing links on queries involving negation
To show that our analysis can be valuable also for different kinds of queries, in Table A.1 we show an analysis for queries involving negation. For such analysis, we split the reasoning tree, defined in Sec. 3, into two subparts, namely positive reasoning tree, composed by the non-negated triples, and the negative reasoning tree, only composed by the negated triples. In particular, in the same spirit of Table 1, by only looking at the positive reasoning tree, we report the percentage of $(q,a)$ pairs that can be reduced to easier types (partial-inference) and the one that cannot be reduced to a simpler type (full-inference). Furthermore, we argue that also the negative reasoning tree contains training triples, but how this propagates to performance is less clear than the positive part, as each method treats negation differently. Table A.1 shows that in both $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, queries can be reduced to easier types, revealing that our analysis also extends to negated queries. The only exceptions are 2in, 2nu1p where the positive reasoning tree consists of a single link, see App. B. Consequently, no reduction to partial inference is possible (see Sec. 3).
Table A.1: Most negated queries have training links in the positive reasoning tree of the QA pairs. Partial-inference and full-inference refer to the positive reasoning tree. ’-’ refers to reductions that are not possible.
$\mathsf{FB15k237}$ $\mathsf{NELL995}$
partial-inference full-inference partial-inference full-inference
2in - 100 - 100
3in 95.4 4.6 93.9 6.1
2pi1pn 98.4 1.6 97.8 2.2
2nu1p - 100 - 100
2in1p 97.4 2.6 95.6 4.4
Table A.2: Even for queries including negation the MRR results are highly influenced by the presence of existing links in the positive reasoning tree of the QA pairs. For example, for 3in queries of $\mathsf{FB15k237}$, the overall MRR of QTO is 16.1, while its MRR on the portion of full-inference QA pairs is only 2.1. Hence, the overall result is highly influence by the vast number of partial inference QA pairs in the benchmark (see Table A.1)
$\mathsf{FB15k237}$ $\mathsf{NELL995}$
Query type method overall partial-inf full-inf overall partial-inf full-inf
2in GNN-QE 6.8 - 6.8 5.5 - 5.5
ULTRAQ 5.3 - 5.3 4.5 - 4.5
CQD 3.3 - 3.3 4.2 - 4.2
CQD-hybrid 4.7 - 4.7 5.1 - 5.1
QTO 10.6 - 10.6 10.2 - 10.2
ConE 5.1 - 5.1 4.6 - 4.6
3in GNN-QE 11.3 11.7 1.1 9.0 9.5 0.6
ULTRAQ 9.6 9.9 0.6 8.2 8.6 0.7
CQD 7.9 8.0 2.7 6.5 6.7 1.1
CQD-hybrid 10.4 10.7 1.1 8.3 8.8 0.7
QTO 16.1 16.7 2.1 14.2 14.9 1.2
ConE 8.0 8.3 1.7 6.5 6.8 1.3
2pi1pn GNN-QE 4.8 4.8 1.8 3.7 3.7 1.5
ULTRAQ 3.6 3.6 1.5 3.2 3.2 1.7
CQD 2.0 2.0 0.6 2.2 2.2 1.4
CQD-hybrid 3.1 3.1 1.1 3.1 3.2 0.8
QTO 8.9 9.0 2.5 7.4 7.5 2.7
ConE 4.1 4.1 1.7 3.0 3.1 1.6
2nu1p GNN-QE 5.0 - 5.0 3.3 - 3.3
ULTRAQ 3.7 - 3.7 2.7 - 2.7
CQD 4.9 - 4.9 4.9 - 4.9
CQD-hybrid 3.2 - 3.2 4.3 - 4.3
QTO 5.3 - 5.3 8.4 - 8.4
ConE 3.3 - 3.3 2.7 - 2.7
2in1p GNN-QE 5.2 5.3 1.7 5.4 5.5 2.7
ULTRAQ 3.8 3.8 1.4 4.7 4.8 1.9
CQD 3.6 3.6 1.4 5.5 5.5 2.6
CQD-hybrid 4.1 4.1 1.5 6.6 6.7 2.5
QTO 7.9 8.1 1.6 7.4 7.6 2.0
ConE 6.3 6.4 3.0 7.5 7.6 4.0
Refer to captionRefer to caption
Figure A.1: Some answers of union queries can be reached by a single missing link, while the other link does not exist as shown for this 2u query and fragments of the KG $\mathsf{FB15k237}$, where $r_{1}=\mathsf{releasedIn}$, $r_{2}=\mathsf{crewMemberOf}$. (Top) Example answer reachable by two missing links. (Bottom) Example answers only reachable by one missing link.
Table A.3: Exploiting the links in the training KG during inference boosts MRR on existing benchmarks, as shown for our CQD-Hybrid model compared to previous SoTA, for several query types. CQD-Hybrid always achieves higher MRR scores when compared to CQD, and outperforms more sophisticated baselines 10/14 times. Best values in bold and second best underlined. For the stratified comparison of CQD-Hybrid with previous SoTA, please refer to the Tables A.5 and A.6.
Method $\mathsf{FB15k237}$ $\mathsf{NELL995}$
2p 3p 2i 3i 1p2i 2i1p 2u1p 2p 3p 2i 3i 1p2i 2i1p 2u1p
GNN-QE 14.7 11.8 38.3 54.1 31.1 18.9 9.7 17.9 15.2 40.0 50.9 29.1 20.5 8.8
ULTRAQ 11.5 8.9 35.7 51.0 29.6 18.6 7.3 11.2 9.7 36.3 47.7 25.1 15.6 8.4
CQD 13.2 7.8 35.0 48.5 27.5 20.7 10.5 22.0 13.4 42.2 51.8 31.5 25.8 16.0
ConE 12.8 11.0 32.6 47.3 25.5 14.0 7.4 16.0 13.8 39.6 50.2 26.1 17.6 11.1
CQD-Hybrid (ours) 15.0 11.0 37.6 52.7 31.2 24.0 10.3 23.8 17.8 44.2 57.8 33.2 28.4 16.6
A.2 Additional comparisons
Table A.6 and Table A.5 show the full results of $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, including the results comparing all methods, all query types, and the subtypes they are reduced to. Furthermore, in Table A.3 we show that our baseline, CQD-hybrid, achieves SoTA performance when compared to non-hybrid baselines, providing evidence that memorization of existing triples leads to inflated performance in the old benchmarks.
Moreover, A.4 show the full results of the union queries, including the old overall results, which include $(q,t)$ pairs having non-existing links in their reasoning tree (see Figure A.1), and the “new overall”, in which those $(q,t)$ pairs are filtered out. In particular, for “2u1p” queries, we observe that for $\mathsf{FB15k237}$, the full-inference queries have a higher MRR than the overall, while for $\mathsf{NELL995}$, we find the MRR results on full-inference $(q,t)$ pairs to be lower than the new overall results as expected. We attribute the results on $\mathsf{FB15k237}$ to the presence of bias in the $(q,t)$ pairs due to their scarcity in the benchmarks. We argue that for the same reason, 2u1p performances are not comparable to 2p. However, we will show that for the full-inference $(q,t)$ pairs in the new benchmark, our claim is confirmed (see Tables D.2, D.3 and D.4).
Moreover, for query types including negation, results in Table A.2 show the same pattern of Table 2, i.e. the MRR of the full-inference QA pairs is much lower than the one of the partial-inference QA pairs and of the overall.
Table A.4: Full inference 2u $(q,t)$ pairs show higher performance than non-filtered, while full-inference 2u1p pairs show lower or comparable performance. Results on union queries for existing benchmarks and comparison between old and new overalls. ‘-’ refers to reductions that are not possible, while ‘/’ to reductions that are possible but that are not present in the data.
$\mathsf{FB15k237}$ $\mathsf{NELL995}$
Query type method overall overall (new) 1p 2u 2u1p overall overall (new) 1p 2u 2u1p
2u GNN-QE 16.2 40.7 / 40.7 - 15.4 34.8 / 34.8 -
ULTRAQ 13.2 33.6 / 33.6 - 10.2 21.3 / 21.3 -
CQD 17.6 32.7 / 32.7 - 19.9 35.9 / 35.9 -
ConE 14.3 29.9 / 29.9 - 14.9 28.2 / 28.2 -
2u1p GNN-QE 13.4 9.7 8.7 50.9 13.1 8.8 8.8 6.3 53.1 2.0
ULTRAQ 10.2 7.3 6.6 33.2 15.0 8.4 7.0 6.6 37.1 6.5
CQD 11.3 10.5 10.4 13.7 14.7 16.0 14.6 14.2 51.8 9.5
ConE 10.6 7.4 7.0 20.3 11.9 11.1 7.0 9.4 43.2 4.0
Table A.5: The best model on all available queries (column “overall”) and the best on the full inference queries (diagonal) is different for 6/7 query types, excluding 1p and 2u, being already of type full-inference. Comparison of MRR scores for SoTA methods for the different subtypes $\mathsf{NELL995}$ queries can be reduced to. Best in bold.
Query type method overall 1p 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p
1p GNN-QE 53.6 53.6 - - - - - - - -
ULTRA-query 38.9 38.9 - - - - - - - -
ConE 60.0 60.0 - - - - - - - -
CQD 60.4 60.4 - - - - - - - -
CQD-hybrid 60.4 60.4 - - - - - - - -
QTO 60.3 60.3 - - - - - - - -
2p GNN-QE 17.9 18.2 6.1 - - - - - - -
ULTRA-query 11.2 11.5 4.6 - - - - - - -
ConE 16.0 16.3 7.2 - - - - - - -
CQD 22.0 22.3 7.6 - - - - - - -
CQD-hybrid 23.8 24.2 6.2 - - - - - - -
QTO 24.7 25.1 7.3 - - - - - - -
3p GNN-QE 15.2 15.1 8.1 3.4 - - - - - -
ULTRA-query 9.7 9.8 5.1 4.1 - - - - - -
ConE 13.8 13.8 8.2 6.5 - - - - - -
CQD 13.4 12.8 7.8 8.5 - - - - - -
CQD-hybrid 17.8 17.2 9.9 8.1 - - - - - -
QTO 22.6 22.3 10.9 7.7 - - - - - -
2i GNN-QE 40.0 41.4 - - 3.6 - - - - -
ULTRA-query 36.3 37.5 - - 3.2 - - - - -
ConE 39.6 40.6 - - 8.0 - - - - -
CQD 42.2 43.3 - - 6.9 - - - - -
CQD-hybrid 44.2 45.7 - - 6.7 - - - - -
QTO 44.2 45.7 - - 5.1 - - - - -
3i GNN-QE 50.9 53.6 - - 10.9 2.2 - - - -
ULTRA-query 47.7 50.0 - - 11.3 1.5 - - - -
ConE 50.2 51.9 - - 18.3 4.8 - - - -
CQD 51.8 53.9 - - 14.4 3.6 - - - -
CQD-hybrid 57.8 61.6 - - 10.7 2.5 - - - -
QTO 56.2 60.0 - - - 2.4 - - - -
1p2i GNN-QE 29.1 42.7 21.4 - 12.0 - 10.2 - - -
ULTRA-query 25.1 39.0 25.6 - 8.2 - 7.0 - - -
ConE 26.1 38.5 25.2 - 10.5 - 8.6 - - -
CQD 31.5 44.0 25.7 - 14.5 - 12.6 - - -
CQD-hybrid 33.2 48.7 24.2 - 13.7 - 9.4 - - -
QTO 32.8 47.4 22.0 - 14.5 - 9.8 - - -
2i1p GNN-QE 20.5 20.8 16.3 - 14.6 - - 20.7 - -
ULTRA-query 15.6 16.5 9.5 - 9.6 - - 11.4 - -
ConE 17.6 17.7 16.4 - 25.7 - - 19.5 - -
CQD 25.8 25.9 21.2 - 26.3 - - 23.6 - -
CQD-hybrid 28.4 28.9 20.4 - 22.8 - - 23.3 - -
QTO 28.2 28.5 20.2 - 24.0 - - 24.4 - -
2u GNN-QE 34.8 - - - - - - - 34.8 -
ULTRA-query 21.3 - - - - - - - 21.3 -
ConE 28.2 - - - - - - - 28.2 -
CQD 35.9 - - - - - - - 35.9 -
CQD-hybrid 35.9 - - - - - - - 35.9 -
QTO 37.6 - - - - - - - 37.6 -
2u1p GNN-QE 8.8 6.3 - - - - - - 53.1 2.0
ULTRA-query 8.4 6.6 - - - - - - 37.1 6.5
ConE 7.0 9.4 - - - - - - 43.2 4.0
CQD 14.6 14.2 - - - - - - 51.8 9.5
CQD-hybrid 15.0 14.5 - - - - - - 53.5 9.6
QTO 16.4 15.8 - - - - - - 61.2 9.6
Table A.6: The best model on all available queries (column “overall”) and the best on the full inference queries (diagonal) is different for 5/7 query types, excluding 1p and 2u, those (q,a) pairs being all of type full-inference. Comparison of MRR scores for SoTA methods for the different subtypes $\mathsf{FB15k237}$ queries can be reduced to. Best in bold.
Query type method overall 1p 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p
1p GNN-QE 42.8 42.8 - - - - - - - -
ULTRA-query 40.6 40.6 - - - - - - - -
ConE 41.8 41.8 - - - - - - - -
CQD 46.7 46.7 - - - - - - - -
CQD-hybrid 46.7 46.7 - - - - - - - -
QTO 46.7 46.7 - - - - - - - -
2p GNN-QE 14.7 14.8 4.7 - - - - - - -
ULTRA-query 11.5 11.5 4.2 - - - - - - -
ConE 12.8 12.8 5.2 - - - - - - -
CQD 13.2 13.3 3.5 - - - - - - -
CQD-hybrid 15.0 15.2 3.5 - - - - - - -
QTO 16.6 16.7 4.0 - - - - - - -
3p GNN-QE 11.8 12.0 4.4 4.8 - - - - - -
ULTRA-query 8.9 9.0 4.6 4.4 - - - - - -
ConE 11.0 11.0 5.4 2.6 - - - - - -
CQD 7.8 7.8 3.4 1.8 - - - - - -
CQD-hybrid 11.0 11.1 3.6 4.4 - - - - - -
QTO 15.6 15.8 4.5 5.0 - - - - - -
2i GNN-QE 38.3 39.3 - - 4.0 - - - - -
ULTRA-query 35.7 36.7 - - 3.4 - - - - -
ConE 32.6 33.3 - - 5.5 - - - - -
CQD 35.0 35.8 - - 7.3 - - - - -
CQD-hybrid 37.6 38.7 - - 3.8 - - - - -
QTO 39.7 40.8 - - 5.7 - - - - -
3i GNN-QE 54.1 56.0 - - 10.9 5.2 - - - -
ULTRA-query 51.0 52.9 - - 9.7 4.3 - - - -
ConE 47.3 48.3 - - 15.6 4.0 - - - -
CQD 48.5 49.6 - - 17.0 6.0 - - - -
CQD-hybrid 52.7 54.8 - - 9.9 4.6 - - - -
QTO 54.6 56.4 - - 15.4 5.4 - - - -
1p2i GNN-QE 31.1 32.8 15.5 - 5.7 - 9.1 - - -
ULTRA-query 29.6 31.5 19.2 - 4.2 - 7.4 - - -
ConE 25.5 26.6 13.9 - 5.4 - 9.7 - - -
CQD 27.5 28.7 13.4 - 7.1 - 9.0 - - -
CQD-hybrid 31.2 33.4 14.8 - 4.8 - 7.0 - - -
QTO 33.8 35.9 15.8 - 6.2 - 7.3 - - -
2i1p GNN-QE 18.9 19.2 8.2 - 6.2 - - 3.4 - -
ULTRA-query 18.6 18.9 8.5 - 5.1 - - 8.1 - -
ConE 14.0 13.8 9.7 - 12.8 - - 5.6 - -
CQD 20.7 21.0 10.5 - 6.7 - - 7.6 - -
CQD-hybrid 24.0 24.6 10.0 - 4.6 - - 7.5 - -
QTO 24.7 25.3 10.3 - 8.6 - - 8.1 - -
2u GNN-QE 40.7 - - - - - - - 40.7 -
ULTRA-query 33.6 - - - - - - - 33.6 -
ConE 29.9 - - - - - - - 29.9 -
CQD 32.7 - - - - - - - 32.7 -
CQD-hybrid 32.7 - - - - - - - 32.7 -
QTO 37.0 - - - - - - - 37.0 -
2u1p GNN-QE 9.7 8.7 - - - - - - 50.9 13.1
ULTRA-query 7.3 6.6 - - - - - - 33.2 15.0
ConE 7.4 7.0 - - - - - - 20.3 11.9
CQD 10.5 10.4 - - - - - - 13.7 14.7
CQD-hybrid 10.3 10.0 - - - - - - 18.5 13.6
QTO 12.2 11.6 - - - - - - 35.3 11.3
Performance analysis of QTO
The stratified MRR performance of QTO (Bai et al., 2023) for the old benchmarks have been included in Tables A.6 and A.5. For a fair comparison, we remark that we used the same link predictor for CQD, CQD-Hybrid, and QTO. Details and hyperameters are available in Sec. F.1. Our analysis reveals that, similarly to the other baselines, QTO performance drops when evaluated on the full-inference $(q,a)$ pairs only (diagonal), w.r.t overall results. Moreover, even when QTO is the SoTA on a certain query type, most of the time it is not so on the portion of full-inference $(q,a)$ pairs only, showing that improving performance on the partial-inference $(q,a)$ pairs not necessarily results on improvements over the full-inference ones.
Moreover, results on $\mathsf{NELL995}$, shown in Table A.5, reveal that CQD-Hybrid outperforms QTO for some query types of $\mathsf{NELL995}$, i.e. 3i, 1p2i, 2i1p, suggesting that even by only setting a score=1 to the training triples it is possible to obtain SoTA results on the old benchmarks. We remark that, while both in CQD-Hybrid and QTO a score=1 is set to the training links, QTO is much more sophisticated than CQD-Hybrid, as they 1) calibrated the scores with a heuristics, 2) store a matrix $|V|\times|V|$ for each relation name containing the score for every possible triples, 3) have a forward/backward mechanism in the reasoning.
A.3 Influence of intermediate existing entities on the results
For query types having intermediate variables, i.e., 2p, 3p, 2i1p, 1p2i, and 2u1p, we analyzed how the number of existing intermediate entities matching the existentially quantified variables could influence the results. We refer to this number as cardinality of existing entities. For example, in Figure 1 the entities “When in Rome” and “Spiderman 2” are existing intermediate entities matching the existential quantified variable $V$ for the query $q_{1}$. The intuition is that while some of these entities simplify the prediction of the answers of the query (see Sec. 3), some others might be misleading for the prediction. Hence, if there is a high number of existing entities in the path, it would be harder for the model to select only the existing entities needed to predict the next link. For this reason, we claim that a high value of cardinality of existing entities would make a $(q,a)$ pair harder.
To support our claim, we analyze the percentage of $(q,t)$ pairs having different values of cardinality, shown in Figure A.3 , and their MRR, using CQD (Arakelyan et al., 2021), shown in Figure A.3. Figure A.3 shows that CQD results are highly influenced by the value of cardinality, having decreasing performance at the increase of the cardinality. Note that by matching this plot with the one in Figure A.3, one can understand if results on a specific query type are highly influenced by a high proportion of $(q,t)$ pairs having the same cardinality. For example, the MRR of 1p2i partial-inference $(q,t)$ pairs in $\mathsf{FB15k237}$ is highly influenced by the presence of about 30% of $(q,t)$ pairs having cardinality of “1”.
Figure A.2: Higher cardinality of intermediate existing entities leads to lower MRR. Influence of the number of existing intermediate variables on the MRR for datasets $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, using CQD.
Figure A.3: Proportion of$(q,t)$ pairs having different cardinality of intermediate existing entities in the old benchmarks.
Table A.7: There is no significant unbalancement in anchor nor relation names for each query type . Most frequent relation name and anchor per query type in the old benchmarks. Highest value in bold.
Query relation name frequency(%) anchor entity frequency(%)
$\mathsf{FB15k237}$
1p film_release_region 7.5 USA 0.7
2p location/location/contains 8.6 USA 4.0
3p location/location/contains 11.6 USA 3.3
2i people/person/profession 16.4 USA 16.4
3i people/person/gender 18.7 USA 30.1
pi people/person/profession 13.7 USA 14.1
ip people/person/profession 8.6 US dollars 5.7
2u film/film_crew_role 14.6 USA 22.5
2u1p taxonomy_entry/taxonomy 9.9 /m/08mbj5d 5.4
2in people/person/profession 15.7 USA 15.2
3in common/webpage/category 20.7 USA 28.5
2pi1pn people/person/profession 12.4 USA 11.7
2nu1p people/person/profession 17.1 USA 14.5
2in1p location/contains 13.4 USA 14.3
$\mathsf{NELL995}$
1p concept:atdate 9.4 concept_stateorprovince_california 0.5
2p concept:mutualproxyfor 8.7 concept_lake_new 1.5
3p concept:mutualproxyfor 15.9 concept_book_new 2.5
2i concept:atdate 22.5 concept_book_new 7.6
3i concept:atdate 21.6 concept_company_pbs 12.0
pi concept:proxyfor 11.4 concept_book_new 6.1
ip concept:subpartof 9.4 concept_athlete_sinorice_moss 3.8
2u concept:atdate 10.8 concept_company_pbs 4.9
2u1p concept:subpartof 11.1 concept_athlete_sinorice_moss 3.3
2in concept:atdate 20.9 concept_book_new 6.8
3in concept:atdate 45.8 concept_book_new 14.2
2pi1pn concept:proxyfor 14.9 concept_book_new 5.7
2nu1p concept:proxyfor 21.9 concept_book_new 8.0
2in1p concept:atdate 15.0 concept_lake_new 6.1
A.4 Imbalances of relation names and anchor nodes
Next, we check if there is some imbalances in both relation names and anchors. Note that if a relation name or an anchor entity is present multiple times in a query, this is counted only once. As shown in Table A.7, for $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ we find that in most cases there is no predominant relation name nor anchor entity. However, there are exceptions, as for example, for anchor entities of 3i queries of $\mathsf{FB15k237}$, where the anchor node “USA” is present in 30.1% of them, and for 2u queries, where it is present in 22.5%. We also notice that there is a predominance of USA as an anchor entity across the vast majority of query types, which is most likely given by the vast presence of this entity in the knowledge graph.
Appendix B Query types including negation
In Figure B.1 we include the query types including the negation operator involved in the analysis described in Sec. A.1. Those queries are obtained by adding a negation to some of the query structures described in Sec. 2. For example, by negating a triple pattern involved in the intersection of a 2i, 3i, 2i1p queries, we obtain respectively:
?$T:(a_{1},r_{1},T)\land\neg(a_{2},r_{2},T),$ (2in)
?$T:(a_{1},r_{1},T)\land(a_{2},r_{2},T)\land\neg(a_{3},r_{3},T),$ (3in)
?$T:\exists V_{1}.((a_{1},r_{1},V_{1})\land\neg(a_{2},r_{2},V_{1})\land(V_{1},r_{3},T)),$ (2in1p)
Moreover, by placing the negation in different triple patterns of the query type 1p2i, we obtain two query types:
?$T:\exists V_{1}.((a_{1},r_{1},V_{1})\land(V_{1},r_{2},T)\land\neg(a_{2},r_{3},T)),$ (2pi1pn)
where the negation is placed in the triple pattern directly involved in the intersection, and
?$T:\exists V_{1}.(\neg((a_{1},r_{1},V_{1})\land(V_{1},r_{2},T))\land(a_{2},r_{3},T)),$ (n2pi1p)
where the negation is placed on the path query 2p. Using De Morgan laws, this query can also be rewritten as:
?$T:\exists V_{1}.(\neg(a_{1},r_{1},V_{1})\lor\neg(V_{1},r_{2},T))\land(a_{2},r_{3},T)),$ (2nu1p)
Note that the visual representation of 2nu1p provided by (Ren & Leskovec, 2020) was misleading, as in practice the negation is applied on the whole 2p query, rather than just on one of its links.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure B.1: Query structures including negation, adapted from Ren & Leskovec (2020), where “2in1p” queries are referred as “inp”. We explicitly mention the number of steps involved in a path or conjunction, as this is a factor of complexity. Analogously, “pin”, “pni” queries from Ren & Leskovec (2020) are now “2pi1np”, and “2nu1p”. See App. B for their logical formulation.
Appendix C More related works
Guu et al. (2015) propose compositional training for embedding methods to predict answers for path queries. GQE (Hamilton et al., 2018) learns a geometric intersection operator to answer conjunctive queries in embedding space; this approach was later extended by Query2Box (Ren et al., 2020), BetaE (Ren & Leskovec, 2020), and GNN-QE (Zhu et al., 2022). FuzzQE (Chen et al., 2022) improves embedding methods with t-norm fuzzy logic, which satisfies the axiomatic system of classical logic. Some recent works such as HypE (Choudhary et al., 2021) and ConE (Zhang et al., 2021) use geometric interpretations of entity and relation embeddings to achieve desired properties for the logical operators. Other solutions to CQA combine neural methods with symbolic algorithms. For example, EmQL (Sun et al., 2020) ensembles an embedding model and a count-min sketch, and is able to find logically entailed answers, while CQD (Arakelyan et al., 2021, 2023) extends a pretrained knowledge graph embedding model to infer answers for complex queries.
Appendix D Additional Analysis on the new benchmarks
D.1 Influence of intermediate existing entities on the results
Similarly to what was done for the old benchmarks in App. A, we analyze the proportion and influence (on the MRR) of the number of intermediate existing entities for CQD, on the new benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$. Figure D.2 show a very similar result to the one in Figure A.3, showing a decreasing MRR at the increase of the value of cardinality for all benchmarks. Moreover, for reference, we also report the proportion of the different $(q,t)$ pairs for each cardinality category in Figure D.2.
Figure D.1: Higher cardinality of intermediate existing entities leads to lower MRR. Influence of the number of existing intermediate variables on the MRR for datasets $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$, using CQD.
Figure D.2: Proportion of$(q,t)$ pairs having different cardinality of intermediate existing entities in new benchmarks.
Query relation name frequency(%) anchor entity frequency(%)
$\mathsf{FB15k237}$+$\mathsf{H}$
1p film_release_region 7.5 USA 0.7
2p /people/person/nationality 15.3 /m/08mbj5d/m/08mbj5d 3.6
3p /people/person/profession 14.2 /m/092ys_y 5.4
4p /dataworld/gardening_hint/split_to 20.0 /m/0dq_5 5.5
2i /people/person/profession 17.2 /m/09c7w0 15.5
3i /people/person/profession 19.6 /m/09c7w0 19.7
4i /people/person/profession 19.2 /m/09c7w0 20.0
pi /people/person/nationality 19.7 /m/09c7w0 14.9
ip /people/person/gender 18.5 /m/048_p 8.9
2in people/person/profession 15.7 USA 15.2
3in /people/person/gender 20.0 /m/09c7w0 20.0
2pi1pn /people/person/nationality 18.4 /m/08mbj5d 13.3
2nu1p people/person/profession 17.1 USA 14.5
2in1p /people/person/profession 13.6 /m/08mbj5d 12.2
2u /people/person/profession 15.6 /m/09c7w0 12.1
2u1p /common/topic/webpage./common/webpage/category 19.4 /m/0kc9f 14.5
$\mathsf{NELL995}$+$\mathsf{H}$
1p concept:atdate 9.4 concept_stateorprovince_california 0.5
2p concept:proxyfor 17.9 concept_sportsteam_ncaa_mens_midwest_regionals 6.4
3p concept:proxyfor 20.0 concept_geopoliticallocation_national 5.9
4p concept:atlocation 20.0 concept_website_new_york_times 6.5
2i concept:atdate 20.0 concept_dateliteral_n2006 5.4
3i concept:atdate 20.0 concept_sportsleague_mlb 5.7
4i concept:atdate 20.0 concept_date_n2004 7.9
pi concept:proxyfor 15.6 concept_book_new 7.8
ip concept:atdate 19.4 concept_lake_new 9.4
2in concept:atdate 20.9 concept_book_new 6.8
3in concept:atdate 20.0 concept_sportsteamposition_center 15.5
2pi1pn concept:proxyfor 20.0 concept_lake_new 15.2
2nu1p concept:proxyfor 21.9 concept_book_new 8.0
2in1p concept:atdate 20.0 concept_lake_new 11.2
2u concept:atdate 20.0 concept_lake_new 2.7
2u1p concept:proxyfor 19.8 concept_lake_new 11.0
$\mathsf{ICEWS18}$+$\mathsf{H}$
1p Make statement 18.8 United States 3.0
2p Make statement 20.0 Russia 6.7
3p Consult 20.0 United States 3.4
4p Make statement 20.0 Yukio Edano 4.4
2i Make statement 20.0 Citizien(India) 7.0
3i Make statement 20.0 Citizien(India) 4.2
4i Make statement 20.0 United States 4.5
pi Make statement 20.0 Unspecified Actor 9.7
ip Make statement 20.0 Russia 6.6
2in Make statement 19.9 Unspecified Actor 17.1
3in Make statement 20.0 Citizien(India) 20.0
2pi1pn Make statement 20.0 United States 11.2
2nu1p Make statement 19.9 Unspecified Actor 16.46
2in1p Consult 20.0 United States 10.5
2u Consult 19.9 China 3.5
2u1p Make statement 19.9 United States 4.9
Table D.1: The new benchmarks are generated such that anchors and relation names cannot appear more than 20% in each query type. Most frequent relation name and anchor per query type in the new benchmarks.
Imbalances of relation names and anchor nodes
When creating the new benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$, to make sure that no anchor entities nor relation name was predominant in each query type, we set a maximum of 20% of $(q,a)$ pairs having the same anchor entity or relation name across all benchmarks and query type, as shown in Table D.1
Stratified analysis on new benchmarks
For reference, we also provide the stratified analysis on the new benchmarks in Tables A.6 and A.5, which shows that 1) the full-inference QA pairs have always a lower MRR than the overall, and 2) the overall MRR is much lower than the 1p-reductions only, as for the new benchmark the proportion of QA pairs is equally distributed.
The same considerations also apply to queries including negation, as shown in Table D.5.
Table D.2: Stratified analysis on $\mathsf{FB15k237}$+$\mathsf{H}$
Query type method overall 1p 2p 3p 4p 2i 3i 4i 1p2i 2i1p 2u 2u1p
1p GNN-QE 42.8 42.8 - - - - - - - - - -
ULTRA-query 40.6 40.6 - - - - - - - - - -
ConE 41.8 41.8 - - - - - - - - - -
CQD 46.7 46.7 - - - - - - - - - -
CQD-hybrid 46.7 46.7 - - - - - - - - - -
QTO 46.7 46.7 - - - - - - - - - -
2p GNN-QE 5.2 11.8 4.4 - - - - - - - - -
ULTRA-query 4.5 9.5 3.9 - - - - - - - - -
ConE 4.6 9.9 4.0 - - - - - - - - -
CQD 4.5 7.4 4.1 - - - - - - - - -
CQD-hybrid 4.8 15.1 3.5 - - - - - - - - -
QTO 4.9 14.4 3.8 - - - - - - - - -
3p GNN-QE 4.0 16.2 4.6 3.6 - - - - - - - -
ULTRA-query 3.5 14.0 3.3 3.2 - - - - - - - -
ConE 3.9 10.1 5.0 3.6 - - - - - - - -
CQD 2.4 3.7 2.2 2.4 - - - - - - - -
CQD-hybrid 3.1 10.7 3.8 2.7 - - - - - - - -
QTO 3.7 21.4 5.8 3.0 - - - - - - - -
4p GNN-QE 4.3 16.1 14.1 4.8 3.8 - - - - - - -
ULTRA-query 3.8 14.4 12.3 4.6 3.3 - - - - - - -
ConE 3.5 12.0 8.8 4.2 3.2 - - - - - - -
CQD 1.1 0.3 0.2 0.7 1.1 - - - - - - -
CQD-hybrid 2.4 9.8 10.5 4.2 2.0 - - - - - - -
QTO 3.9 18.2 21.7 5.5 3.2 - - - - - - -
2i GNN-QE 6.0 20.4 - - 5.4 - - - - - - -
ULTRA-query 5.2 20.7 - - 4.6 - - - - - - -
ConE 9.1 15.5 - - 8.8 - - - - - - -
CQD 11.3 18.4 - - 10.8 - - - - - - -
CQD-hybrid 6.0 22.0 - - 5.4 - - - - - - -
QTO 8.7 24.4 - - 8.0 - - - - - - -
3i GNN-QE 8.8 32.2 - - - 2.5 7.5 - - - - -
ULTRA-query 7.2 32.4 - - - 2.4 5.9 - - - - -
ConE 10.3 25.0 - - - 6.9 9.0 - - - - -
CQD 12.8 30.5 - - - 8.5 11.5 - - - - -
CQD-hybrid 8.6 38.2 - - - 2.7 7.1 - - - - -
QTO 10.1 40.0 - - - 6.5 8.2 - - - - -
4i GNN-QE 19.6 47.9 - - - 18.3 16.5 15.8 - - - -
ULTRA-query 16.4 48.9 - - - 17.5 13.2 12.9 - - - -
ConE 20.3 36.9 - - - 24.9 18.2 15.8 - - - -
CQD 23.8 45.1 - - - 27.0 20.9 19.8 - - - -
CQD-hybrid 17.4 58.3 - - - 17.2 13.2 14.6 - - - -
QTO 20.2 60.2 - - - 25.1 16.3 15.6 - - - -
1p2i GNN-QE 5.6 28.1 6.3 - - 3.0 - - 3.5 - - -
ULTRA-query 5.3 31.4 7.6 - - 1.9 - - 3.1 - - -
ConE 3.8 16.7 5.1 - - 2.5 - - 2.6 - - -
CQD 6.0 20.1 7.2 - - 3.5 - - 4.2 - - -
CQD-hybrid 5.5 33.2 7.7 - - 2.6 - - 3.0 - - -
QTO 6.1 32.0 8.1 - - 3.8 - - 3.6 - - -
2i1p GNN-QE 9.9 13.5 6.1 - - 16.8 - - - 7.2 - -
ULTRA-query 10.1 13.8 5.7 - - 18.4 - - - 7.2 - -
ConE 7.9 10.8 4.8 - - 10.4 - - - 7.2 - -
CQD 11.5 12.2 6.4 - - 21.7 - - - 10.5 - -
CQD-hybrid 12.9 23.0 6.0 - - 24.7 - - - 8.2 - -
QTO 13.5 23.5 6.6 - - 27.3 - - - 9.0 - -
2u GNN-QE 32.4 - - - - - - - - - 32.4 -
ULTRA-query 29.4 - - - - - - - - - 29.4 -
ConE 22.8 - - - - - - - - - 22.8 -
CQD 40.0 - - - - - - - - - 40.0 -
CQD-hybrid 42.2 - - - - - - - - - 42.2 -
QTO 30.6 - - - - - - - - - 30.6 -
2u1p GNN-QE 10.0 7.6 - - - - - - - - 54.0 6.5
ULTRA-query 8.3 7.4 - - - - - - - - 42.2 4.8
ConE 6.0 6.0 - - - - - - - - 17.1 4.7
CQD 10.6 9.2 - - - - - - - - 13.9 8.4
CQD-hybrid 12.0 10.5 - - - - - - - - 44.0 8.1
QTO 11.2 10.0 - - - - - - - - 53.3 6.7
Table D.3: Stratified analysis on $\mathsf{NELL995}$+$\mathsf{H}$
Query type method overall 1p 2p 3p 4p 2i 3i 4i 1p2i 2i1p 2u 2u1p
1p GNN-QE 53.6 53.6 - - - - - - - - - -
ULTRA-query 38.9 38.9 - - - - - - - - - -
ConE 60.0 60.0 - - - - - - - - - -
CQD 60.4 60.4 - - - - - - - - - -
CQD-hybrid 60.4 60.4 - - - - - - - - - -
QTO 60.3 60.3 - - - - - - - - - -
2p GNN-QE 8.0 21.6 6.3 - - - - - - - - -
ULTRA-query 6.1 16.2 5.0 - - - - - - - - -
ConE 7.9 15.3 7.1 - - - - - - - - -
CQD 9.6 21.7 8.0 - - - - - - - - -
CQD-hybrid 9.0 28.4 6.6 - - - - - - - - -
QTO 9.8 25.9 7.7 - - - - - - - - -
3p GNN-QE 6.0 33.3 14.1 3.9 - - - - - - - -
ULTRA-query 4.1 14.7 9.4 3.1 - - - - - - - -
ConE 6.7 21.0 12.2 5.5 - - - - - - - -
CQD 4.2 6.2 5.6 3.7 - - - - - - - -
CQD-hybrid 6.1 22.2 10.9 4.4 - - - - - - - -
QTO 8.0 42.1 15.3 5.2 - - - - - - - -
4p GNN-QE 4.7 4.9 25.0 11.9 2.8 - - - - - - -
ULTRA-query 4.2 30.5 17.1 11.5 2.9 - - - - - - -
ConE 5.2 27.3 17.3 7.5 4.5 - - - - - - -
CQD 2.0 2.7 1.4 0.9 2.0 - - - - - - -
CQD-hybrid 3.5 34.9 19.3 8.1 2.1 - - - - - - -
QTO 7.0 54.4 29.1 12.4 4.9 - - - - - - -
2i GNN-QE 10.7 24.5 - - 9.8 - - - - - - -
ULTRA-query 7.9 18.7 - - 7.3 - - - - - - -
ConE 21.8 23.3 - - 21.1 - - - - - - -
CQD 18.5 26.3 - - 17.6 - - - - - - -
CQD-hybrid 12.1 28.7 - - 11.0 - - - - - - -
QTO 14.6 30.5 - - 13.5 - - - - - - -
3i GNN-QE 13.3 34.7 - - - 11.7 10.6 - - - - -
ULTRA-query 10.2 26.7 - - - 11.4 7.1 - - - - -
ConE 23.6 31.8 - - - 22.5 21.5 - - - - -
CQD 19.6 34.6 - - - 17.4 17.1 - - - - -
CQD-hybrid 14.4 43.3 - - - 11.7 11.8 - - - - -
QTO 15.8 46.2 - - - 14.1 12.8 - - - - -
4i GNN-QE 19.4 51.9 - - - 24.6 14.3 13.5 - - - -
ULTRA-query 15.6 42.6 - - - 24.9 12.2 8.2 - - - -
ConE 27.6 47.9 - - - 35.0 21.8 23.8 - - - -
CQD 24.8 52.9 - - - 30.6 17.7 20.8 - - - -
CQD-hybrid 20.4 65.1 - - - 24.9 14.2 14.5 - - - -
QTO 20.9 68.3 - - - 26.8 13.8 15.4 - - - -
1p2i GNN-QE 16.0 55.9 19.5 - - 9.1 - - 7.3 - - -
ULTRA-query 15.8 55.0 23.0 - - 5.5 - - 5.9 - - -
ConE 14.9 32.6 18.9 - - 10.7 - - 7.0 - - -
CQD 18.9 54.3 22.1 - - 12.0 - - 10.0 - - -
CQD-hybrid 17.4 64.7 19.0 - - 12.4 - - 8.1 - - -
QTO 17.6 58.4 20.6 - - 13.4 - - 7.9 - - -
2i1p GNN-QE 13.5 26.2 12.9 - - 17.0 - - - 8.2 - -
ULTRA-query 9.3 22.2 7.6 - - 22.3 - - - 6.0 - -
ConE 11.8 22.6 11.1 - - 23.2 - - - 8.4 - -
CQD 22.6 29.9 20.1 - - 36.9 - - - 16.1 - -
CQD-hybrid 21.2 35.8 18.9 - - 34.5 - - - 13.7 - -
QTO 21.1 36.3 19.2 - - 36.0 - - - 13.4 - -
2u GNN-QE 47.5 - - - - - - - - - 47.5 -
ULTRA-query 28.1 - - - - - - - - - 28.1 -
ConE 39.9 - - - - - - - - - 39.9 -
CQD 46.3 - - - - - - - - - 46.3 -
CQD-hybrid 46.4 - - - - - - - - - 46.4 -
QTO 49.1 - - - - - - - - - 49.1 -
2u1p GNN-QE 9.8 11.4 - - - - - - - - 42.9 5.2
ULTRA-query 9.5 12.3 - - - - - - - - 36.9 5.1
ConE 8.8 10.3 - - - - - - - - 21.9 6.3
CQD 18.5 18.8 - - - - - - - - 53.5 13.4
CQD-hybrid 19.3 20.5 - - - - - - - - 63.7 13.1
QTO 18.9 21.1 - - - - - - - - 58.2 11.7
Table D.4: Stratified analysis on $\mathsf{ICEWS18}$+$\mathsf{H}$
Query type method overall 1p 2p 3p 4p 2i 3i 4i 1p2i 2i1p 2u 2u1p
1p GNN-QE 12.2 12.2 - - - - - - - - - -
ULTRA-query 6.3 6.3 - - - - - - - - - -
ConE 3.5 3.5 - - - - - - - - - -
CQD 16.6 16.6 - - - - - - - - - -
CQD-hybrid 16.6 16.6 - - - - - - - - - -
QTO 16.6 16.6 - - - - - - - - - -
2p GNN-QE 0.9 2.2 0.8 - - - - - - - - -
ULTRA-query 1.2 2.1 1.1 - - - - - - - - -
ConE 0.9 1.9 0.9 - - - - - - - - -
CQD 2.6 5.4 2.4 - - - - - - - - -
CQD-hybrid 2.6 10.0 2.0 - - - - - - - - -
QTO 2.6 8.4 2.2 - - - - - - - - -
3p GNN-QE 0.5 1.0 1.2 0.4 - - - - - - - -
ULTRA-query 1.2 2.1 1.5 1.2 - - - - - - - -
ConE 0.9 1.5 1.4 0.9 - - - - - - - -
CQD 1.5 2.6 3.2 1.4 - - - - - - - -
CQD-hybrid 1.5 9.8 4.4 1.2 - - - - - - - -
QTO 1.4 6.2 3.7 1.2 - - - - - - - -
4p GNN-QE 0.4 1.6 1.4 1.0 0.3 - - - - - - -
ULTRA-query 0.8 2.9 1.9 0.9 0.7 - - - - - - -
ConE 0.6 1.8 1.8 0.9 0.6 - - - - - - -
CQD 1.1 1.6 1.5 1.4 1.1 - - - - - - -
CQD-hybrid 0.9 31.4 39.3 20.1 0.7 - - - - - - -
QTO 1.1 14.3 5.4 2.5 0.7 - - - - - - -
2i GNN-QE 16.1 30.0 - - 2.4 - - - - - - -
ULTRA-query 7.0 12.3 - - 2.1 - - - - - - -
ConE 1.2 1.9 - - 0.8 - - - - - - -
CQD 13.0 22.1 - - 4.3 - - - - - - -
CQD-hybrid 15.0 27.3 - - 3.3 - - - - - - -
QTO 15.7 28.3 - - 3.4 - - - - - - -
3i GNN-QE 26.5 58.5 - - - 19.3 2.0 - - - - -
ULTRA-query 11.7 25.9 - - - 8.0 1.6 - - - - -
ConE 0.5 1.2 - - - 0.6 0.2 - - - - -
CQD 19.5 42.3 - - - 13.4 3.2 - - - - -
CQD-hybrid 25.6 58.5 - - - 15.8 2.7 - - - - -
QTO 24.7 56.7 - - - 16.3 2.5 - - - - -
4i GNN-QE 34.0 75.0 - - - 44.9 13.9 1.6 - - - -
ULTRA-query 15.9 37.1 - - - 19.1 6.0 1.0 - - - -
ConE 0.3 0.9 - - - 0.4 0.2 0.1 - - - -
CQD 24.0 54.2 - - - 28.8 10.2 2.3 - - - -
CQD-hybrid 33.2 78.9 - - - 39.4 11.8 2.1 - - - -
QTO 31.5 73.6 - - - 38.0 11.9 18.6 - - - -
1p2i GNN-QE 19.1 62.5 30.1 - - 2.5 - - 2.3 - - -
ULTRA-query 8.8 19.6 13.2 - - 28 - - 2.2 - - -
ConE 1.2 2.6 1.5 - - 1.0 - - 0.8 - - -
CQD 17.1 38.5 25.8 - - 50.1 - - 4.4 - - -
CQD-hybrid 17.5 66.9 26.2 - - 4.1 - - 3.5 - - -
QTO 18.4 59.3 28.0 - - 4.4 - - 3.6 - - -
2i1p GNN-QE 3.5 4.5 3.0 - - 5.7 - - - 3.4 - -
ULTRA-query 1.3 2.3 1.2 - - 1.6 - - - 1.2 - -
ConE 1.6 2.4 1.6 - - 2.4 - - - 1.3 - -
CQD 6.9 6.1 6.0 - - 11.4 - - - 6.8 - -
CQD-hybrid 5.8 9.2 4.6 - - 14.8 - - - 5.0 - -
QTO 6.2 10.3 5.5 - - 13.1 - - - 5.1 - -
2u GNN-QE 7.6 - - - - - - - - - 7.6 -
ULTRA-query 3.3 - - - - - - - - - 3.3 -
ConE 1.1 - - - - - - - - - 1.1 -
CQD 6.8 - - - - - - - - - 6.8 -
CQD-hybrid 6.8 - - - - - - - - - 6.8 -
QTO 6.7 - - - - - - - - - 6.7 -
2u1p GNN-QE 1.1 0.8 - - - - - - - - 4.7 0.8
ULTRA-query 1.2 1.5 - - - - - - - - 2.0 1.0
ConE 0.9 1.0 - - - - - - - - 1.3 0.8
CQD 5.9 3.2 - - - - - - - - 11.6 5.6
CQD-hybrid 5.6 2.7 - - - - - - - - 17.8 4.5
QTO 4.9 3.3 - - - - - - - - 17.8 3.7
Table D.5: Also for queries involving negation, there is no clear SoTA method for the new benchmarks. The MRR on the new benchmarks is significantly lower than the old ones. For example, for 3in queries on $\mathsf{FB15k237}$+$\mathsf{H}$, QTO has an MRR of only 3.1, while for $\mathsf{FB15k237}$ it was 16.1.
$\mathsf{FB15k237}$+$\mathsf{H}$ $\mathsf{NELL995}$+$\mathsf{H}$ $\mathsf{ICEWS18}$+$\mathsf{H}$
Query type method ovr part-inf full-inf ovr par-inf full-inf ovr par-inf full-inf
2in GNN-QE 6.8 - 6.8 5.5 - 5.5 4.5 - 4.5
ULTRAQ 5.3 - 5.3 4.5 - 4.5 2.3 - 2.3
CQD 3.3 - 3.3 4.2 - 4.2 1.5 - 1.5
CQD-hybrid 4.7 - 4.7 5.1 - 5.1 1.7 - 1.7
QTO 10.6 - 10.6 10.2 - 10.2 4.9 - 4.9
ConE 5.1 - 5.1 4.6 - 4.6 1.7 - 1.7
3in GNN-QE 6.5 9.0 1.0 6.4 7.2 1.0 6.9 12.6 1.4
ULTRAQ 5.5 7.5 0.6 5.9 6.2 0.9 4.8 7.1 2.4
CQD 2.6 5.9 2.5 1.5 6.0 1.3 2.9 4.0 1.6
CQD-hybrid 1.6 8.4 1.2 1.2 7.9 0.9 4.0 6.5 1.3
QTO 3.1 13.9 2.3 2.3 11.8 1.8 8.7 14.5 2.8
ConE 4.9 6.4 2.0 6.0 5.8 2.8 2.9 3.3 2.0
2pi1pn GNN-QE 3.7 4.0 1.4 5.8 7.0 1.6 0.9 1.0 0.5
ULTRAQ 2.6 2.8 1.3 4.3 5.2 1.7 1.2 1.3 0.9
CQD 0.6 1.3 0.5 1.5 3.2 1.4 0.2 0.7 0.1
CQD-hybrid 1.0 2.7 0.9 1.4 5.9 1.1 0.3 1.0 0.2
QTO 2.0 7.3 1.7 3.1 9.4 2.6 1.2 3.4 1.0
ConE 2.9 2.7 1.5 3.7 4.9 2.0 1.1 1.3 1.0
2nu1p GNN-QE 5.0 - 5.0 3.3 - 3.3 3.5 - 3.5
ULTRAQ 3.7 - 3.7 2.7 - 2.7 2.2 - 2.2
CQD 4.9 - 4.9 4.9 - 4.9 2.7 - 2.7
CQD-hybrid 3.2 - 3.2 4.3 - 4.3 2.2 - 2.2
QTO 5.3 - 5.3 8.4 - 8.4 3.0 - 3.0
ConE 3.3 - 3.3 2.7 - 2.7 0.9 - 0.9
2in1p GNN-QE 3.3 2.4 1.4 4.4 4.6 2.1 0.8 1.0 0.5
ULTRAQ 2.2 2.2 1.2 3.6 3.7 1.6 1.6 2.2 1.4
CQD 1.2 1.3 1.2 2.6 3.5 2.6 0.9 0.2 0.9
CQD-hybrid 1.3 1.1 1.2 2.4 4.1 2.3 1.1 1.9 1.0
QTO 1.5 5.4 1.3 2.4 7.9 1.9 0.9 2.5 0.9
ConE 3.6 2.8 2.2 6.4 7.0 3.9 1.3 2.3 1.2
Appendix E Queries with four hops
Logical formulation of the 4p and 4i queries we introduced in the benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$
?$T:$ $\ \exists V_{1},V_{2},V_{3}.(a_{1},r_{1},V_{1})\land(V_{1},r_{2},V_{2})\land$ (4p)
$\ (V_{2},r_{3},V_{3})\land(V_{3},r_{4},T),$
?$T:$ $\ (a_{1},r_{1},T)\land(a_{2},r_{2},T)\land$ (4i)
$\ (a_{3},r_{3},T)\land(a_{4},r_{4},T).$
Appendix F Hyperparameters
In this section, we detail the hyperparameters used for each dataset and model. Old and new benchmarks, the generation scripts, and the implementation of CQD-hybrid are included as supplemental materials in a GitHub repo. [^6] https://github.com/april-tools/is-cqa-complex/tree/main
F.1 Old benchmarks
GNN-QE
We did not tune hyperparameters, but re-used the ones provided in the official repo. [^7] https://github.com/DeepGraphLearning/GNN-QE/tree/master/config
ULTRA
We did not tune hyperparameters, but re-used the ones provided in the official repo. [^8] https://github.com/DeepGraphLearning/ULTRA/tree/main/config/ultraquery
CQD
As mentioned in Sec. 5, we re-used the pre-trained link predictor (Trouillon et al., 2017) provided by the authors. [^9] https://github.com/Blidge/KGReasoning/ However, we tuned CQD-specific hyperparameters, namely the CQD beam “k”, ranging from [2,512] and the t-norm type being “prod” or “min” In Table F.1 we provide the hyperparameter selection for the old benchmarks $\mathsf{FB15k237}$ and $\mathsf{NELL995}$. Also note that we normalize scores with min-max normalization.
Dataset 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p
k tn k tn k tn k tn k tn k tn k tn k tn
$\mathsf{FB15k237}$ 512 p 8 p 128 p 128 p 256 p 64 p 512 m 512 m
$\mathsf{NELL995}$ 512 p 2 p 2 p 2 p 256 p 256 p 512 m 512 m
Table F.1: CQD hyperparameters old benchmarks. tn = tnorm. p=prod, m=min. Note that for 1p neither the k nor the tnorm are needed.
ConE
We did not tune hyperparameters but re-used the ones provided in the official repo. [^10] https://github.com/MIRALab-USTC/QE-ConE/blob/main/scripts.sh
CQD-Hybrid
For CQD-Hybrid, to make the comparison with CQD fair, we re-used the hyperparameters found for CQD and fixed an upper bound value for the CQD beam “k” to 512, even when there are more existing entities matching the existentially quantified variables, to match the upper bound of the “k” used for CQD. Additionally, the scores of the pre-trained link predictor are normalized between $[0,0.9]$ using min-max normalization, and a score of $1$ is assigned to the existing triples.
F.2 New benchmarks
For $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$, we re-used the same models trained for the old benchmarks, and using the same hyperparameters presented in Sec. F.1. Instead, for the new benchmark $\mathsf{ICEWS18}$+$\mathsf{H}$ we trained every model. The used hyperparameters for each model are presented in the following:
GNN-QE
For GNN-QE, we tuned the following hyperparameters: (1) batchsize, with values 8 or 48, and concat hidden being True or False, while the rest are the same used for the old benchmarks and do not change across benchmarks. For $\mathsf{ICEWS18}$+$\mathsf{H}$ the best hyperparameters are “batchsize=48” and “concat hidden=True”.
ULTRAQ
Being a zero-shot neural link predictor, we re-used the same checkpoint provided in the official repo, as for F.1.
CQD
For $\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{FB15k237}$+$\mathsf{H}$ we re-used the pre-trained link-predictor and the same hyperparameters found for the old benchmarks F.1. Instead, for the newly created $\mathsf{ICEWS18}$+$\mathsf{H}$, we train ComplEx (Trouillon et al., 2017) link predictor with hyperparameters “regweight” 0.1 or 0.01, and batch size 1000 or 2000, with the best being, respectively 0.1 and 1000. Moreover, in F.2 are shown the hyperparameters for CQD on the new benchmark $\mathsf{ICEWS18}$+$\mathsf{H}$.
Dataset 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p 4i 4p
k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn
$\mathsf{ICEWS18}$+$\mathsf{H}$ 32 p 2 p 2 p 2 p 512 p 256 p 2 m 2 m 2 p 2 p
Table F.2: CQD hyperparameters new $\mathsf{ICEWS18}$+$\mathsf{H}$ benchmark. tn = tnorm. p=prod, m=min
ConE
We re-used the same hyperparameters of $\mathsf{NELL995}$ for $\mathsf{ICEWS18}$+$\mathsf{H}$.
Appendix G Statistics
G.1 KG splits statistics
The statistics, i.e., number of entities, relation names, training/validation/test links, of the knowledge graphs used in this paper are shown in Table G.1.
Table G.1: Statistics of knowledge graphs used to generate complex queries
Dataset Entities Relation Names Training Links Validation Links Test Links Total Links
$\mathsf{FB15k237}$ 14,505 237 272,115 17,526 20,438 310,079
$\mathsf{NELL995}$ 63,361 200 114,213 14,324 14,267 142,804
$\mathsf{ICEWS18}$+$\mathsf{CQ}$ (new) 20,840 250 213,304 25,048 24,689 263,041
The KG splits of $\mathsf{NELL995}$ and $\mathsf{FB15k237}$ are the same of the one used in (Ren & Leskovec, 2020; Ren et al., 2020).
G.2 New benchmark statistics
G.2.1 $\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{FB15k237}$+$\mathsf{H}$
Training queries
were not changed from the one proposed by Ren & Leskovec (2020); Ren et al. (2020).
Validation and Test
queries are generated to ensure a balanced distribution of hardness across the sub-query types to which a given query can be reduced. As described in Sec. 6, we generate queries such that each sub-query type contains exactly 10,000 $(q,t)$ pairs, along with an additional 10,000 pairs corresponding to full-inference queries (i.e., non-reducible cases). To achieve this, we adapt the algorithm from (Ren et al., 2020), sampling queries until the required number of $(q,t)$ pairs for each sub-query type is reached. For instance, for 2p queries, we continue sampling until we obtain 10,000 $(q,t)$ pairs reducible to 1p and 10,000 full-inference queries. When a query is sampled, all its corresponding answers are included in the benchmark until the target of 10,000 $(q,t)$ pairs per subtype is met. If incorporating all answers would exceed this limit, a random subset is selected to ensure exactly 10,000 pairs. Similarly, once a sub-query type reaches the 10,000-pair threshold, any additional answers belonging to that category are discarded.
Only exceptions are 1p, 2in, and 2nu1p, which being already of full-inference type we kept the same as in (Ren & Leskovec, 2020; Ren et al., 2020).
G.2.2 ICEWS18+H
Training queries
were generated using the same strategy used in (Ren & Leskovec, 2020; Ren et al., 2020) , namely by (1) creating 1p queries by considering every triple in the training set, (2) retrieve the number $n$ of generated 1p queries, (3) generate $n$ queries for all training query types.
Validation and test queries
were generated using the same strategy detailed in Sec. G.2.1. Moreover, since we generated this benchmark from scratch, we also had to generate the 1p, 2in, 2nu1p queries, which following (Ren & Leskovec, 2020; Ren et al., 2020), were generated for each triple in the validation and test split.